Two cars are 500 miles apart. Both start at the same time and travel toward one another. They meet 5 hours later. If the speed of one car is 20 mph faster than the other, what is the speed of each car? Must show work.
Answer by rossiv53(27)

let x = the rate of first car and x+20 =the rate of the second car.
The distance the first car travels is the product of 5x since it traveled 5 hours. The second is 5(x+20). The total distance they travel is 500 miles.
5x+5(x+20)=500
5x+5x+100=500
10x+100=500
10x+100-100=500-100
10x=400
x=40
So, 1st car's speed is 40 and since the second car travels 20 mph faster, its speed is 60 mph.
Good luck!